will save labor costs by as much as $35,000 per You are year over 10 years. The equipment costs S200,000. Should you purchase the equipment if your interest rate is 12%?
the eq $100 $100 $100 $100 $100 OA V4 I$100 (P/A, i, 6)-$100 (P/F, i, 4) (FIP, ,4) O B. ¼·S 100 (FAL 3) + $100 (PAL 2) O c-V-($100 (FA i, 6) , $100 (FRI 2)] (PF, i, 2) OD. ¼=$100 (FA,i,4)-$100 + $100 (PAL2)
You have $15,000 available for looking for? in stock. You are looking for a growth stock whose value can grow to $40,000 over five years. What kind of growth rate are The annual growth rate you are looking for is (Round to two decimal places)
Answer #1

(Question 1)

Net present worth (NPW) of the equipment ($) = - 200,000 + 35,000 x P/A(12%, 10) = - 200,000 + 35,000 x 5.6502**

= - 200,000 + 197,757

= - 2,243

Since NPW is negative, I will not purchase this equipment.

**From P/A Factor table

NOTE: As per Answering Policy, 1st question is answered.